How many atoms of zinc are in 1.00 g


1
Expert's answer
2022-02-23T02:03:02-0500

Moles of zinc "=\frac{1.00}{65.38}=0.0153moles"

One mole of zinc has "6.022\times 10^{23}atoms"

Atoms of zinc in 1.00g "=0.0153\times 6.022\times 10^{23}=9.21366\times 10^{21}atoms"